# An Experimental Study on the Rashomon Effect of Balancing Methods in Imbalanced Classification

### v1

@@ 5 Conclusions
In this paper, the effect of balancing methods on predictive multiplicity was investigated by addressing different RQs. It is important because exploring the Rashomon set provides some advantages and challenges. It can be utilized to select models that satisfy additional properties without compromising accuracy, such as fairness, interpretability, and stability [32]. However, it may be problematic because the models in the Rashomon set rely on different variables and the blind selection of the model can produce more profitable or unfavorable predictions to the individuals [33].
@@
We conduct experiments on benchmark datasets with varying imbalanced ratios. For the RQ1, we found that balancing methods increase Rashomon metrics. In the RQ2, we investigated the effect of balancing methods with the newly proposed metric we proposed on the model behavior through variable importance order and did not observe a statistically significant order change that contradicts the findings of Alarab and Prakoonwit [6]. However, considering that they conducted their experiments on only 2 data sets, we think the results we obtained on a larger number of datasets are more comprehensive and reliable. The results we obtained above that the balancing methods inflate the Rashomon metrics making it necessary to develop solutions. First, in the RQ3, we examined whether the partial resampling approach suggested by Kamalov et al. [9] could be a solution. However, we found that partial resampling does not work. Then, by expanding the performance-gain plot and monitoring the balance between performance and Rashomon metrics, we showed that choosing the method that best suits the data could be a solution. With its easy-to-use structure, the performance-gain plot is quite flexible and expandable across the intended Rashomon metrics.
@@
In conclusion, we would like to draw attention again to the recent discussions of the use of balancing methods with an additional dimension. Our experiments show that balancing methods increase the predictive multiplicity and increase the risks involved in the process of model selection from the Rashomon set. Therefore, researchers should consider the risk of multiplicity when using balancing methods, which are a haven in imbalanced classification problems. The use of the extended performance gain plot for the Rashomon effect is important to monitor this issue and conduct the modeling process responsibly.
6 Further Research
@@
There are still opportunities for further exploration in this field. To mitigate the predictive multiplicity, identifying the source of the Rashomon effect is a hot topic [33]. First, Komorniczak et al. [34] stated in their study that oversampling methods increase data complexity. The question of whether there is a connection between data complexity and the Rashomon effect can be investigated, e.g., ”Does the increase in data complexity due to oversampling methods lead to an increase in the Rashomon effect?”. Second, similar to the study by Junior and Pisani [35], it can be investigated whether cost-sensitive methods, such as resampling methods, have a lesser impact on the Rashomon effect. Third, special cases of imbalance situations such as class overlap, and small disjunct, as mentioned by authors such as Garcia et al. [36] and Prati et al. [37], can also be examined.
Supplemental Materials
@@
The materials for reproducing the experiments performed and the benchmark datasets are given in the repository: https://github.com/mcavs/ECML2024_Imbalanced_Rashomon_Paper.

### v3

@@ 5 Conclusions
This paper investigated the effect of one commonly used data-centric AI approach in the training data development phase, balancing methods on predictive multiplicity by addressing different RQs. It is essential because exploring the Rashomon set provides some advantages and challenges. It can be utilized to select models that satisfy additional properties without compromising accuracy, such as fairness, interpretability, and stability [39]. However, it may be problematic because the models in the Rashomon set rely on different variables, and the blind selection of the model can produce more profitable or unfavorable predictions to the individuals [40].
@@
We conduct experiments on benchmark datasets with varying imbalanced ratios. For the RQ1. How do the balancing methods affect the predictive multiplicity of the models in the Rashomon set? We found that balancing methods increase Rashomon metrics. In the RQ2. How do the balancing methods affect the variable importance order discrepancy of the models in the Rashomon set? We investigated the effect of balancing methods with a new metric we proposed on the model behavior through variable importance order and did not observe a statistically significant order change that contradicts the findings of Alarab and Prakoonwit [9]. However, considering that they conducted their experiments on only two data sets, we think the results we obtained on a larger number of datasets are more comprehensive and reliable. The results we obtained above show that the balancing methods inflate the Rashomon metrics, making it necessary to develop solutions. First, in the RQ3. Can partial resampling be a solution to the model behavior change? We examined whether the partial resampling approach suggested by Kamalov et al. [14] could be a solution. However, we found that partial resampling does not work. Then, focusing on RQ4. Can the extended version of the performance gain plot be a solution to monitor the trade-off between performance gain and multiplicity? and expanding the performance-gain plot and monitoring the balance between performance and Rashomon metrics, we showed that choosing the method that best suits the data could be a solution. Its easy-to-use structure makes the performance-gain plot flexible and expandable across the intended Rashomon metrics.
@@
In conclusion, it is vital to consider the recent discussions on balancing methods during the training data development phase of data-centric AI, incorporating an additional dimension. Our experiments show that balancing methods increase the predictive multiplicity and the risks involved in the model selection process from the Rashomon set. Therefore, researchers should consider the risk of multiplicity when using balancing methods, which are a haven in imbalanced classification problems. Using the extended performance-gain plot for the Rashomon effect is essential to monitor this issue and conduct the modeling process responsibly.
Further Research
@@
There are still opportunities for further exploration in this field. To mitigate the predictive multiplicity, identifying the source of the Rashomon effect is a hot topic [40]. First, it is stated that oversampling methods increase data complexity [41]. The question of whether there is a connection between data complexity and the Rashomon effect can be investigated, e.g., “Does the increase in data complexity due to oversampling methods lead to an increase in the Rashomon effect?”. Second, similar to the study in [42], it can be investigated whether cost-sensitive methods, such as resampling methods, have a lesser impact on the Rashomon effect. Third, special cases of imbalance situations, such as class overlap and small disjunct, as mentioned in [43, 44], can also be examined.
Acknowledgment.
This study is funded by the Scientific and Technological Research Council of Turkiye (TÜBİTAK) through the 2224-A Grant Program for Participation in Scientific Meetings Abroad (application no. 1919B022401776) and Eskisehir Technical University Scientific Research Projects Commission under grant no. 24ADP116.
Supplemental Materials.
@@ Acknowledgment.
The materials for reproducing the experiments and the benchmark datasets are available in the repository: https://github.com/mcavs/ECML2024_Imbalanced_Rashomon_Paper.